Tierney Company begins operations on April 1. Information from job cost sheets shows the following. Manufacturing Costs Assigned Job Month Number April MayJune Completed 10 $6,800 $5,000 11 6,000 5,300 $ 2.600 June 12 13 14 May 2,300 April 5,300 5,600 June 6,600 4,600 Not complete Job 12 was completed in April. Job 10 was completed in May. Jobs 11 and 13 were completed in June. Each job was sold for 25% above its cost in the month following completion. What is the balance in Work in Process Inventory at the end of each month? Work in Process Inventory April 30 $ May 31$ June 30 $ Attempts: 0 of 7 used C Check AnswerWhat is the balance in Work in Process Inventory at the end of each month? Work in Process Inventory April 30 May 31 June 30 LINK TO TEXT LINK TO TEXI VIDEO: SIMILAR EXERCISE What is the balance in Finished Goods Inventory at the end of each month? Finished Goods Inventory April 30 May 31 June 30 What is the gross profit for May, June, and July? Gross Profit May une uly Click if you would like to Show Work for this question: pen Show Work
Answer #1
a
Work in Process Inventory
April 30 12800 =6800+6000
May 31 23200 =6000+5300+5300+6600
June 30 11200 =6600+4600
b
Finished goods Inventory
April 30 2300
May 31 11800 =6800+5000
June 30 24800 =6000+5300+2600+5300+5600
c
Gross Profit
May 575 =2300*0.25
June 2950 =11800*25%
July 6200 =24800*25%